OpenStudy (anonymous):

What is the LCD of 5/12 and -9/16 ?

OpenStudy (welshfella):

you can find the LCD of 2 numbers from their prime factors 12 = 2*2*3 16 = 2*2*2*2

OpenStudy (welshfella):

where you have duplicated factors - here we have 2 pairs of '2' we only include one of the duplicates so the lcd = 2*2*2*2*3 = 48 (we include all the other numbers)

OpenStudy (anonymous):

cool ok so is the answer 48

OpenStudy (welshfella):

yes - do you understand the reasoning?

OpenStudy (anonymous):

yes very will thank you nobody was help me today

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

what

OpenStudy (welshfella):

thats means you are welcome

OpenStudy (anonymous):

ohh np

OpenStudy (welshfella):

another example: LCD of 6 and 9 6 = 2*3 9 = 3*3 - only include one of the duplicate 3's so its 2*3*3 = 18

OpenStudy (anonymous):

uhh cool math is strange so times

OpenStudy (welshfella):

and if you have say 3 numbers then you include only 1 of a triplicate 6, 9 and 12 6 = 2*3 9 = 3*3 12 = 2*2*3 there is a pair of 2's and a trio of 3's so you only include 1 of these 3's and 1 of the 2's

OpenStudy (anonymous):

What is the LCD of 5/6 and -3/8 ?

OpenStudy (anonymous):

24

OpenStudy (welshfella):

6 = 2*3 8 = 2*2*2 what do you think?

OpenStudy (welshfella):

yes:- 2*2*2*3

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

good luck

OpenStudy (anonymous):

thanx